Newtown Corp. has to choose between two mutually exclusive projects. If it chooses project A, Newtown Corp. will have the opportunity to make a similar investment in three years. However, if it chooses project B, it will not have the opportunity to make a second investment. The following table lists the cash flows for these projects. If the firm uses the replacement chain (common life) approach, what will be the difference between the net present value (NPV) of project A and project B, assuming that both projects have a weighted average cost of capital of 13%?

Cash Flow

Project A Project B
Year 0: –$17,500 Year 0: –$40,000
Year 1: $10,000 Year 1: $8,000
Year 2: $16,000 Year 2: $16,000
Year 3: $15,000 Year 3: $15,000
Year 4: $12,000
Year 5: $11,000
Year 6: $10,000

* $13,626

* $12,023

* $17,633

* $10,420

* $16,030

Newtown Corp. is considering a five-year project that has a weighted average cost of capital of 14% and a NPV of $80,720. Newtown Corp. can replicate this project indefinitely. What is the equivalent annual annuity (EAA) for this project?

* $23,512

* $27,039

* $24,688

* $29,390

* $21,161
Answer #1

Answer 1:

Correct answer is:

* $16,030

Explanation:

If the firm uses the replacement chain (common life) approaches the cash flows and NPV will be as follows:

Difference in NPV = 24169 - 8139 = $16,030

Hence option E is correct and other options A, B, C and D are incorrect.

Answer 2:

Correct answer is:

* $23,512

Explanation:

PV factor of annuity for 5 years at 14% discount = (1 - 1 / (1 + 14%) 5 ) / 14% = 3.433081

Equivalent annual annuity (EAA) = NPV / PV factor of annuity = $80,720 / 3.433081 = $23,512

Hence option A is correct and other options B, C, D, and E are incorrect.